### v1

# Designing and Evaluating Hint Generation Systems for Science Education

Abstract.
Large language models are influencing the education landscape, with students relying on them in their learning process. Often implemented using general-purpose models, these systems are likely to give away the answers, which could hinder conceptual understanding and critical thinking. We study the role of automatic hint generation as a pedagogical strategy to promote active engagement with the learning content, while guiding learners toward the answers. Focusing on scientific topics at the secondary education level, we explore the potential of large language models to generate chains of hints that scaffold learners without revealing answers. We compare two distinct hinting strategies: static hints, pre-generated for each problem, and dynamic hints, adapted to learners’ progress. Through a quantitative study with 41 participants, we uncover different preferences among learners with respect to hinting strategies, and identify the limitations of automatic evaluation metrics to capture them. Our findings highlight key design considerations for future research on hint generation and intelligent tutoring systems that seek to develop learner-centered educational technologies.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
Large language models have reshaped how students interact with digital learning tools. These models are increasingly used for everyday study practices, from gaining conceptual understanding to supporting problem-solving and writing (Levine et al., 2025; Rahma and Fithriani, 2024; Syarifah and Fakhruddin, 2024). At the same time, online learning continues to expand at unprecedented scales. According to UNESCO’s Global Education Monitoring Report (2023) (Antoninis et al., 2023), the number of students enrolled in massive open online courses reached at least 220 million in 2021, reflecting both the demand for and the accessibility of scalable educational resources. While prior work in education research has established a strong correlation between the student-teacher ratio and learning outcomes (Koc and Celik, 2015), teacher recruitment and retention crisis remains a worldwide conundrum (Huang et al., 2022; Williams III et al., 2022; Castro, 2022). ”Even where general teacher supply and demand are in balance, many countries face shortages of specialist teachers and shortages in schools serving disadvantaged or isolated communities.(Schleicher, 2012). These constraints create a growing reliance on automated support to bridge the gap in access to personalized instruction.
Despite their rapid adoption, many educational technologies are deployed with limited evaluation before reaching learners, and most evidence of the impact of education technology comes from the richest countries. Antoninis et al. (2023) reports that in the United Kingdom, only 7% of education technology companies had conducted randomized controlled trials, while a survey of teachers and administrators in 17 states within the United States revealed that only 11% requested peer-reviewed evidence prior to adoption. This context underscores the need for research that not only develops new forms of automated support, but also studies their effectiveness, limitations, and design trade-offs.
Recent advances in LLM reasoning have spurred exploration of their potential for education, continuing a trajectory that began well before LLMs with intelligent tutoring systems (ITS) (Chu et al., 2025; Alhafni et al., 2024; Yan et al., 2024; Abd-Alrazaq et al., 2023). However, as in the pre-LLM era, the majority of research has centered on structured domains such as programming and mathematics, where correctness can be precisely defined and evaluation is straightforward. In contrast, relatively little attention has been paid to unstructured problem spaces, such as science learning, where solutions may be expressed in varied ways. To address this gap, our work focuses on hint generation systems for scientific problem solving, specifically designing hints for factoid questions as the first step towards scaffolding learners to enhance understanding and critical thinking in science domains (see Figure 1).
Figure 1. Illustration of automatic chain-of-hint generation for scientific question answering.
We introduce the concept of a chain-of-hints, where information is revealed incrementally to scaffold the learner’s reasoning process. Inspired by the principles of human tutoring, this approach mirrors how teachers provide just-in-time interventions — offering minimal guidance first, and providing progressively more explicit support as needed (Wood et al., 1976). We systematically compare two strategies for generating and delivering these hints: pre-generated static hints and adaptive dynamic hints that respond to learner’s progress. Through this lens, we investigate how learners perceive the helpfulness of different hinting strategies, and derive design principles that can guide the development of pedagogically effective educational systems. We ground our study in the following research questions:
  • RQ1: What characteristics make hints useful to learners?
  • RQ2: How do learners perceive adaptive (dynamic) versus pre-generated (static) hinting strategies?
  • RQ3: To what extent do automatic evaluation metrics align with learners’ perceptions of hint quality?
In order to answer these questions, we conducted an evaluation study with 41 participants. The findings revealed the promise and complexity in designing effective hint generation systems. Learners valued hints that clarified concepts, provided additional context and incremental support. Yet, their preferences diverged over certain characteristics of static and dynamic hints, especially hints that took a context-agnostic approach to guide learners towards the answers. There is thus an implicit trade-off between short-term satisfaction and long-term accumulation of cognitive debt. Finally, we find that the automatic evaluation metrics are only partially aligned with the learner perceptions of hints’ helpfulness.
Building on these insights, we discuss design implications for future hint generation and tutoring systems, applicable beyond factoid question answering. We call for learner-centered design practices that identify and account for diverse needs and preferences, rather than using general-purpose models. We also advocate for tutoring frameworks that integrate self-regulated learning with gradually faded support, enabling learners to develop metacognitive awareness and inhibiting over-reliance on automated assistance. In doing so, we position hint generation as a pathway to tackle the rising problem of cognitive debt due to over-reliance on readily available AI assistants designed to provide complete responses. This work makes three key contributions through our exploration of scientific hints as a form of interactive learning support: (1) we present results from an evaluation study with 41 learners toward understanding the components that make hints useful; (2) a suite of metrics for automatic evaluation of chain-of-hints across five dimensions; and (3) design implications for the development and evaluation of intelligent tutoring systems.
2. Related Work
2.1. Automatic Hint Generation: Past and Present
As a core component of intelligent tutoring systems, automatic hint generation has been widely studied across educational domains, with sustained interest in how hints can improve students’ learning outcomes and reasoning skills. Prior work can be broadly categorized into three paradigms based on their underlying framework to provide hints: (i) rule-based models, (ii) data-driven models, and (iii) LLM-based models (Stamper et al., 2024).
Rule-based models formed the backbone of early intelligent tutoring systems. These systems relied heavily on expert-designed learner models, which allowed them to provide effective, problem-specific feedback (Mitrovic et al., 2001; Aleven et al., 2009). However, their dependence on expert input made them difficult to scale and limited their generalizability across diverse learners and domains.
Data-driven models offered a more scalable alternative, particularly in domains with well-structured problems such as programming (Barnes and Stamper, 2008; Rivers and Koedinger, 2017; Obermüller et al., 2021; Jin et al., 2012; Zimmerman and Rupakheti, 2015; Paaßen et al., 2018; Price et al., 2016; Rolim et al., 2017). These models typically followed a deterministic framework consisting of three components: (1) a corpus of diverse candidate solutions, (2) a matching algorithm to align learner attempts with the closest example, and (3) a path traversal algorithm to guide learners toward the correct solution. A rich body of surveys documents the innovations in this space (Le et al., 2013; Crow et al., 2018; McBroom et al., 2021; Mahdaoui et al., 2022). While effective in programming education, data-driven models were constrained by their dependence on high-quality training data, their limited adaptability beyond formal logic domains, and their difficulty in handling solution diversity.
To overcome these limitations, recent work has increasingly explored LLM-based hint generation models. Large language models could offer various advantages, including improved reasoning ability, adaptability, and domain generalizability (Bommasani, 2021; Naveed et al., 2025; Kasneci et al., 2023; Minaee et al., 2024). While programming remains a popular testbed for LLM-based systems (Rogers et al., 2025; Demirtaş et al., 2025; Barros et al., 2025; Borchers and Shou, 2025; Ma et al., 2024; Liang et al., 2024; Kazemitabaar et al., 2024), the scope has expanded to other subjects with structured solution space, like mathematics. For instance, Wang et al. (2023b) curated a benchmark (ReMath) co-developed with teachers to evaluate step-by-step math tutoring, with the aim to use LLMs to assist math tutors in remediating student mistakes. Pal Chowdhury et al. (2024) proposed MWPTutor, a framework for generating math word problem hints by reformulating questions around the learner’s next step. Pardos et al. (2023) introduced OATutor, an open-source ITS platform to support adaptive tutoring research across multiple subjects. Building on this, Pardos and Bhandari (2024) evaluated ChatGPT-generated hints within OATutor, comparing them to human tutor hints across algebra and statistics. Their results demonstrate statistically significant learning gains from ChatGPT-generated hints.
Recent work has also examined automatic feedback in more open-ended contexts. For instance, Lieb and Goel (2024) developed NewtBot, a LLM-based tutoring chatbot to support secondary school students in physics in German. They explored several pedagogy-grounded and context-grounded prompting strategies against a general purpose model, however, their findings reveal no statistically significant difference across the prompting strategies, likely due to small group size. In design education, Duan et al. (2024) utilized LLMs to generate heuristic feedback on UI mockups, offering constructive but sometimes inconsistent suggestions. In legal education, Weber et al. (2024) developed LegalWriter, an intelligent writing support system that can help novice students produce more structured and persuasive case solutions. These efforts highlight that while early work concentrated on structured topics, researchers are beginning to explore how LLM-based feedback can support learning in more open-ended domains.
Despite this growing diversity, we identify two key reasons why research on hint generation has focused most heavily on structured domains. First, evaluation is more straightforward: solutions adhere to strict rules, allowing researchers to measure correctness with high reliability. Second, these domains align with the broader aim to improve formal reasoning capabilities of LLMs, developing benchmarks for establishing state-of-the-art performance (Chen et al., 2021b; Zhuo et al., 2024; Jiang et al., 2024; Paul et al., 2024; Glazer et al., 2024; Mishra et al., 2022; Li et al., 2020; Lu et al., 2023). By contrast, subjects involving a natural language solution space are underexplored within the intelligent tutoring systems community, due in part to the difficulty of evaluation. Through this work, we aim to contribute toward the development of hint generation methods that extend beyond formal logic settings.
2.2. Automatic Evaluation of Hints
Evaluating the quality of the generated hints at scale requires reliable automatic metrics. Prior work has proposed two families of methods: reference-based and reference-free evaluation.
Reference-based metrics assume access to expert-written hints and compare generated hints against them. For example, Price et al. (2019) proposed QualityScore, which evaluates programming hints by measuring similarity to tutor-authored hints using abstract syntax tree (AST) representations (McCarthy, 1964; Knuth, 1968). While such methods can provide fine-grained comparisons, they are inherently limited by the availability and coverage of expert references—often infeasible to scale across domains.
Reference-free metrics, in contrast, aim to predict hint helpfulness without requiring gold-standard references. These methods either leverage learner interaction data or adopt learner-agnostic success measures. Interaction-based metrics quantify how a learner’s state changes after receiving a hint. For instance, Paaßen et al. (2018) computed error over two distances: (i) between the predicted post-hint state and the true next state, and (ii) between the predicted post-hint state and the learner’s final state. Learner-agnostic approaches instead assess hints through generalizable outcomes such as correctness or efficiency. For programming hints, Rivers and Koedinger (2017) evaluated hints based on whether they eventually reached a correct solution and the length of the hint sequence. More recently, for natural language hints, Mozafari et al. (2025) proposed several metrics such as relevance, readability, convergence, familiarity, and answer leakage to capture a hint’s quality in open-ended domains.
In our work, we extended this line of research by investigating a broader set of pedagogically-grounded evaluation metrics for a chain of hints that better reflect the sequential nature of hint delivery. We also evaluated the alignment of these metrics to learners’ needs and experiences, identifying the blind spots of automatic evaluation. This analysis surfaces gaps in current metrics, and highlights open challenges for designing metrics that better capture hint utility from a learner-centered perspective.
Table 1. Dataset statistics for the automatic evaluation and human evaluation splits.
Dataset Split #instances #instances per Subject Average #words
Question Answer Context
Automatic Evaluation 1000
Biology (518), Chemistry (160),
Geology (143), Physics (162),
Miscellaneous (17)
13.23 1.52 66.89
Human Evaluation 30
Biology (8), Chemistry (7),
Geology (7), Physics (8)
13.20 1.27 61.30
2.3. Learner-centered evaluation of generated hints
Human-centered evaluation is indispensable for building effective hint generation systems. Because learners vary in their needs, pace, and strategies, evaluations must consider how different designs support diverse populations. Prior work has explored a range of evaluation settings, shaped by study design choices and deployment targets. We organize these into four broad approaches: i) third-party annotation, ii) randomized controlled trials, iii) classroom-scale deployments, and iv) longitudinal studies.
Third-party annotation. To assess the pedagogical quality of LLM-generated responses, Tack and Piech (2022) conducted a pairwise human-in-the-loop comparison study, with annotators rating responses along dimensions such as talking like teachers, understanding students, and helping improve student understanding. Their findings revealed that while LLMs often match teachers in conversational uptake, they lag significantly in their ability to foster deeper understanding.
Randomized controlled trials (RCTs). RCTs are among the most common designs in prior work, allowing direct comparisons between feedback strategies in controlled settings. For instance, Lieb and Goel (2024) evaluated feedback strategies for secondary physics, while Duan et al. (2024) examined feedback for UI mockup design. Both highlight how targeted experimental setups provide causal evidence of effectiveness, though typically over shorter time frames.
Classroom-scale deployments. A number of studies have examined hinting systems in authentic classroom contexts, capturing the social and ecological dynamics of learning. For example, Kazemitabaar et al. (2024) deployed CodeAid, a programming assistant designed to provide hints without revealing full solutions, while Rogers et al. (2025) studied programming support in a classroom setting through MatlabTutee, a tool designed with an emphasis on the learning-by-teaching principle.
Longitudinal studies. Longer-term deployments allow researchers to capture sustained patterns of tool use and their influence on learner development, especially within the nascent stages of prototype development. Rogers et al. (2025), for instance, distinguished between high-usage and low-usage groups in a two month-long deployment, finding that heavy users of their proposed system (MatlabTutee) reported the highest self-efficacy gains, underscoring the value of extended exposure in evaluating learning technologies.
Across these evaluation modes, researchers emphasized different outcome dimensions depending on the domain and artifact design. Helpfulness or usefulness consistently appears as a central measure of perceived utility. Other dimensions included self-efficacy (Rogers et al., 2025), attitudes toward AI assistance (measured via instruments such as the godspeed questionnaire (Bartneck et al., 2009) or short version of the user experience questionnaire (Schrepp et al., 2014)), enjoyability and cognitive load (Lieb and Goel, 2024), and task-specific factors such as analogical reasoning in the case of bio-inspired design training (Chen et al., 2024b). These variations underscore how human evaluations illuminate distinct facets of educational support that automatic metrics alone cannot capture.
In our work, we adopted a within-subjects randomized controlled trial to evaluate hints for scientific question answering. Our study emphasized on hint’s usefulness, insightfulness, and oversimplification, while also examining how these perceptions align or diverge from automatic evaluation metrics.
3. Scientific Hint Generation
Expert tutors are able to provide high quality support to students to help them meet their learning goals. These tutors take into consideration each individual’s learning objectives (Hoque, 2016; Sönmez, 2017) and prior knowledge (Orsmond and Merry, 2011; Biggs, 1996) to provide meticulously crafted direct instructions (Kozloff et al., 1999; Kim and Axelrod, 2005; Rosenshine, 2008) to guide the learners to the correct answer, molding them into deep conceptual thinkers (Rillero, 2016) with an enhanced sense of accomplishment (VanLehn, 2011; Merrill et al., 1992; Arroyo et al., 2014). Through this work, we aim to explore the capabilities of state-of-the-art reasoning models to provide this scaffolded guidance to learners to reach the correct answer. Limiting the focus to scientific hint generation, in this section, we formally define the chain-of-hints generation task (Section 3.1), and describe the dataset we use for the evaluation (Section 3.2) of several LLM-based baselines (Section 3.3). We also briefly describe the automatic evaluation metrics (Section 3.4) we use to select the model for user study, and provide the detailed description and analysis in Appendix A.
3.1. Problem Definition
Drawing from the just-in-time incremental interventions portrayed by expert human tutors, we formally define the chain-of-hints generation task as follows: given a question $q$ and answer $a$, a hint generation system $\mathbf{H}$ generates a sequence of hints $\textbf{h}=\langle h_{1},h_{2},...,h_{k}\rangle$, such that:
[…]
where $P(a|.)$ denotes the probability of a user successfully answering the question. We propose two variants of the task depending on whether the hint generation system takes prior attempts into account. In the static setting, the hint generation function is defined as
$\mathbf{H}:\{q,a,\langle h_{1},h_{2}\ldots h_{i-1}\rangle\}\mapsto h_{i}$
where the next hint $h_{i}$ only depends on the question, the correct answer and the sequence of previously provided hints. In the dynamic setting, the function additionally conditions on the learner’s past incorrect attempts, i.e.,
$\mathbf{H}:\{q,a,\langle h_{1},h_{2},\ldots h_{i-1}\rangle,\langle a_{1}^{\times},a_{2}^{\times}\ldots\rangle\}\mapsto h_{i}$
Figure 2. Performance of best models from five baseline model families evaluated using automatic evaluation metrics. Orange (solid) bars denote the static hint generation results and blue (striped) bars denote the dynamic hint generation results.
3.2. Dataset
To evaluate the ability of large language models to generate hints, we use the SciQ dataset (Welbl et al., 2017a) for our evaluation. SciQ dataset is a science exam question answering dataset comprising of 13,679 multiple choice questions spread across biology, chemistry, physics and geology; spanning across elementary level to college introductory level questions. Since having multiple choices for each question hinders the ability to test the quality of generated hints, we discard the distractor choices, and view the task as a free-form short answer question answering task. We use two different dataset splits for the automatic and human evaluation as follows (see Table 1 for detailed statistics) -
  • Dataset for automatic evaluation To extensively evaluate the hint generation capabilities of the LLM-based baselines, we use the 1,000 examples from the validation split of the SciQ dataset.
  • Dataset for human evaluation As the original SciQ dataset comprised of multiple-choice question answers, we hand-pick 30 questions from the test split of SciQ dataset to ensure valid open-ended questions with unambiguous answer, while ensuring equal representation of all four subjects in the final quiz.
3.3. Baselines
To extensively evaluate the capabilities of large language models to generate static and dynamic hints, we explore a diverse range of open-source model families. We limit our exploration on smaller open-source language models for three reasons - 1) open-source language model weights are constant, unlike API-based models, 2) there are privacy concerns with using closed-source models, and 3) any real-world application for learning feedback will have computational constraints, and thus we do not explore large language models that require multiple GPUs to execute. More specifically, in this work, we explore the following language models for the scientific hint generation task - DeepSeek-R1 (1.5B, 7B, 8B, 14B, 32B) (Guo et al., 2025), Gemma3 (1B, 4B, 12B, 27B) (Team et al., 2025), Mistral-Small (24B) [^1] https://huggingface.co/mistralai/Mistral-Small-24B-Instruct-2501, Phi4 (14B) (Abdin et al., 2024), and Qwen3 (0.6B, 1.7B, 4B, 8B, 14B, 30B, 32B) (Yang et al., 2025). We use an in-context learning approach to generate the static and dynamic chain-of-hints (refer to Figures 10 and 11 for the static and dynamic hint generation prompts respectively).
3.4. Automatic Evaluation Metrics
Reliable automatic evaluation metrics are critical in advancing research on automatic hint generation, and intelligent tutoring systems more broadly (Mousavinasab et al., 2021; Jangra et al., 2024). These automatic evaluation metrics provide a fast feedback cycle in early-stage system development. At the same time, they serve as a safeguard against premature deployment of underdeveloped systems. Exposing students to low-quality hint generation systems may not only undermine the learning outcomes, but also jeopardize learners’ trust in educational technologies more broadly. Having robust automatic evaluation metrics also helps ensure comparability and reproducibility of research findings across different studies. To automatically evaluate the pedagogical quality of the chain-of-hints in ways that align with learner needs and goals, we discuss several automatic evaluation metrics, either by adapting metrics from prior works, or proposing new ones wherever necessary. We briefly describe the evaluation metrics here, and provide detailed definitions in Appendix A.1.
We investigate a suite of evaluation metrics spanning five dimensions: information gain, redundancy, consistency, readability, and leakage. To capture the degree to which hints help learners progress toward correct answers, we propose information gain ($InfoGain$), which leverages a language model to compare question-answering performance with and without hints. We implement two variants: $InfoGain_{mean}$, the average gain across all hints in a chain, and $InfoGain_{comb}$, the maximum gain achieved by the entire hint chain.
Drawing on variation theory (Ling Lo, 2012), which emphasizes the value of presenting learners with diverse perspectives, we assess redundancy by adapting self-referenced redundancy from summarization research (Chen et al., 2021a), with the goal of minimizing repetitive information. To ensure factual correctness, we introduce consistency, computed using AlignScore (Zha et al., 2023), a metric designed to evaluate factual alignment between generated text and the source material.
To prevent hints from oversimplifying tasks or directly revealing answers, we implement leakage with two complementary measures: a string-matching approach ($leakage_{EM}$) and a LLM-based approach ($leakage_{LLM}$) designed to detect subtle forms of answer disclosures. Finally, to evaluate clarity, we apply established readability measures, including the Flesch–Kincaid grade level (Kincaid et al., 1975), Flesch reading ease (Flesch, 1979), and Dale–Chall readability score (Dale and Chall, 1948).
For model selection for our human evaluation study, we combine information gain, redundancy, consistency, and leakage into a single aggregate score to holistically evaluate the content of the generated hints.
Experimental Setup. We evaluate all baselines over 1,000 instances from the validation split of the SciQ question answering benchmark (Welbl et al., 2017b). We explore two different hint generation settings to quantify the hint generation capabilities of these baselines - static and dynamic. For the static setting, we prompt the LLM baseline to generate 4 hints given a question answer pair. For the dynamic setting, we used the three distractor choices from the SciQ multiple choice questions as the incorrect responses, and interleaved them within each hint to replicate a learner interaction for the in-context learning prompt. We provide the prompts for static and dynamic hint generation in Figures 10 and 11 respectively.
Figure 2 summarizes the performance of the best-performing baselines from each model family. Comparing static and dynamic hint generation, we find that dynamic hints generally exhibit higher redundancy and lower consistency scores. This reduced content diversity, combined with elevated leakage rates, results in lower aggregate scores for dynamic hint chains relative to static hints. Across both settings, Mistral-Small-24b achieves the highest overall performance. Given the large number of baselines and evaluation settings, we report detailed experimental results in Appendix A.2. There, we examine trends across metrics, contrasts between static and dynamic strategies, and the impact of model parameters on hint quality. Later, in Section 5.4, we analyze the effectiveness of these automatic evaluation metrics to evaluate the quality of hints grounded on learner feedback from our user study.
Selecting a Model for User Study: We select Mistral-Small-24b model to conduct the human evaluation study for two main reasons - i) it achieved the highest aggregate score in both static and dynamic hint generation settings, and ii) due to it’s great instruction following capabilities, the hints generated from Mistral-Small-24b followed our desired structure in the output with minimal post-processing, an important property for the human evaluation study with online hint generation for the dynamic hint generation setting.
4. User Study Design
We conducted a within-subjects human evaluation study with 41 participants to identify the characteristics of high-quality hints that learners might find most useful (RQ#1). We asked the participants to engage with our user interface in an unmoderated setting at their own convenience to obtain feedback in a more realistic self-study environment preventing any researcher bias.
Figure 3. Our quiz interface for human evaluation study.
Procedure. Our study comprised of three parts: pre-quiz briefing, quiz, and post-quiz briefing. In the pre-quiz assessment phase, participants completed a preliminary questionnaire collecting demographic information (gender, ethnicity, age, highest educational attainment) and self-reported familiarity with four science domains: physics, chemistry, biology, and earth science (measured on a 5-point scale from ‘novice’ to ‘expert’).
In the quiz assessment phase, the participants where given 30 free-form short answer questions standardized for all participants (8 physics, 8 biology, 7 chemistry and 7 geology questions). The quiz consisted of three sections, each containing 10 questions. With a within-subjects study design in mind, the first section served as the control condition with no hints provided, while the 2nd and 3rd sections featured either static or dynamic hints, assigned randomly to prevent order bias or recency bias in post-quiz briefing survey. For each question, participants had the option to request up to four hints before submitting an answer, and were allowed five attempts before their response was marked incorrect (see Figure 3). After each question, participants were presented with the correct answer and asked to evaluate the hints they had received. To assess the quality of each hint, we focused on three dimensions: participants’ overall satisfaction (five-point Likert scale), whether the hint provided new knowledge (informativeness), and whether it directly disclosed the correct answer (leakage). This structured feedback process allowed us to systematically assess participants’ perceptions of hint utility and informativeness, and assess the quality of our automatic evaluation metrics. After each section, participants were asked to answer four questions: one assessing the difficulty of the quiz on a five-point Likert scale, one rating the overall quality of hints in previous section on a five-point Likert scale, and two open-ended questions about the positives and negatives of hints shown in previous section. Due to lack of hints in the first Section, we only asked the first question in the 1st section’s survey. Participants progressed through the quiz at their own pace and had the option to take a break after completing each section, allowing them to pause and resume the study as needed.
Upon completing the quiz questions and the associated surveys, participants responded to a short post-quiz briefing survey, where they were informed about the two different hint strategies used in the 2nd and 3rd sections of the quiz. Their entire quiz interaction, including questions, answers, attempted answers and hints shown was displayed for reference, and based on their experience and perception, we asked three questions - one asking to select the most helpful hint strategy, one asking to select the hint strategy that helped improve the question understanding, and an open-ended question about their observed differences in hints between the two strategies (dynamic vs static). We also provided space for participants to share their thoughts about the study at the end of this post-quiz briefing survey. The study was reviewed and approved by the IRB at our institution.
Figure 4. Statistics of 41 human evaluation study participants.
Type Count
Age 24.81 years (mean), 3.85 years (std)
Ethnicity
Asian (29), Multi-racial (3), White (3),
Black (2), Middle-eastern (2), Undisclosed (2)
Education Background Masters (19), Bachelors (10), PhD (9), Undisclosed (2), High school (1)
Gender Female (23), Male (16), Non-binary (1), Undisclosed (1)
Table 2. Demographics of 41 study participants.
Participants. We recruited 46 participants by circulating our study through university mailing lists and slack channels across several universities in North America. We solicited participants who had studied science subjects in high school, lived in the United States, and were at least 18 years old. Each participant received a 20 dollars gift card for their participation. We do not include in our analysis the data from five participants that asked less than five hints to prevent introducing noise from learners who did not meaningfully engage with the system. From the remaining 41 participants in our study, we obtained feedback over 1,149 hints asked (639 static, 510 dynamic). 20 participants were first shown static hints and then dynamic hints, while for the other 21 participants the order was reversed to prevent ordering bias. On average, participants took 28.44 minutes to complete the study, where they asked 28.02 hints and successfully answered 21.85/30 questions. We provide the distribution of these statistics in Figure 4, and the aggregate participant demographics in Table 2. We redact all identifiable details from the data, and use participant IDs when quoting them in our findings.
Analysis. To identify the effectiveness of a hint in answering questions, we conducted extensive quantitative analysis over three self-reported quality indicators: 1) participant satisfaction rating, an explicit measure of hint quality obtained directly from the learners on a 5-point Likert scale, 2) insightfulness, the participant’s response to the question, ”Did you learn something new from the hint?”, and 3) answer disclosure, the participant’s response to the question, ”Did the hint give away the answer?”. We also analyzed the influence of multiple hints, and explore the preference of different hinting strategies across a sequence of hints amongst the participants. To compare effectiveness across the two different hint generation strategies (static vs dynamic), we used an aggregate of these fine-grained indicators in tandem with the participant-reported quality assessment scores in post-section survey. We also analyzed the responses to the open-ended questions using inductive analysis to identify patterns in participants’ feedback from post-section and post-quiz surveys (Thomas, 2003).
5. Findings
In this section, we present findings from both qualitative and quantitative analyses. We begin with RQ#1, examining what makes hints useful to learners. This includes both the value of a single hint (Section 5.1) and the added complexity of chains of hints (Section 5.2). For RQ#2, we compare how learners perceive different hinting strategies, contrasting static and dynamic approaches (Section 5.3). Finally, for RQ#3, we address the alignment between automatic evaluation metrics and learners’ own judgments of hint quality (Section 5.4).
5.1. When a Single Hint Helps (or Hurts)
Our analysis reveals that at a sentence-level, hints portrayed several key characteristics of a successful feedback such as clarity and simplicity, yet recurring issues emerged around answer leakage and ambiguity. While most characteristics were clearly divided into good and bad across participants, we found that the preference was divided among the participants for certain types of hints. Our inductive analysis of learners’ feedback highlights several themes around perceived hint quality:
Improved conceptual understanding. Overall, participants marked 26.63% (306/1149) hints as informative, with a moderate correlation between the self-reported satisfaction and informativeness ($corr_{pearson}=0.4955$), and several participants explicitly noted that hints helped improve their understanding of the question, not just leading them to the answer. As P-22 described, ”hints encouraged to make associations and engage in reflection from a scientific perspective”. From out inductive analysis of qualitative responses, we identified three distinct ways in which hints supported this process. New perspectives: some hints reframed the questions to open new lines of thoughts. Linking to prior knowledge: other hints helped connect the new concepts to their own knowledge, helping them form associations. Supplementary knowledge: finally, several hints helped enrich participants’ knowledge with additional facts. Figure 5 (top left example) illustrates an interaction where hints help improve the participant’s understanding about the question by indicating a geological activity that involves magma rising to the surface to lead them to the correct answer volcanoes.
The impact of domain familiarity on hint helpfulness. Participants described the usefulness of hints to be closely tied to whether the concepts referenced were already familiar to them. If the participants were familiar with the concept, then hints were often helpful in reaching the answer, especially when hints drew on colloquial expressions or widely known facts. For instance, for question ”What is the most abundant metal of the earth’s crust?”, the hint ”Consider metals that are commonly used in everyday items like foil and cans.” was very effective in aiding the participants to answer the question. At the same time, the hints proved less effective when the concepts themselves were unfamiliar. In those cases, hints did not provide sufficient grounding context, leaving participants uncertain about how to proceed. Participants even reported feeling reluctant to request hints in such cases, as P-21 reflected: ”Some questions I don’t know about the words, it doesn’t make sense to take hints cuz i still don’t know the words…”. Using the number of attempted answers as a proxy for engagement (and to some extent familiarity), we observed similar trends, higher engagement instances yielding higher satisfaction ratings (see Figure 6). Therefore, while familiarity could amplify the helpfulness of a hint by triggering recognition and recall, it could also hinder the learning, as without prior knowledge to anchor to, the same hints risked compounding confusion rather than resolving it.
In contrast, several participants emphasized that hints were unhelpful when they already recognized the concept but could not recall the precise term. As P-31 explained, ”If i remember the answer but it was on the tip of my tongue and cant remember the exact one, the hints do not help at all.”. This highlights a broader constraint of our short-form factoid question-answering setup: the system requires participants to supply the exact target term, meaning that hint effectiveness was often judged not only by whether it supported conceptual understanding, but also by whether it guided the learner toward the expected answer [^2] While we adopted a fuzzy match for our underlying answer assessment implementation, using in-context learning with an LLM (we used OpenAI’s gpt-3.5-turbo for the answer assessment task.), there were instances where it expected the exact lexical form (e.g., for the correct answer hydroxide ions, the submission OH- was marked incorrect). To ensure the efficacy of our assessment prompt for a functioning study, we test out the quality of the assessment model from a previous pilot study with 697 answers over the same 30 questions. While assessment model (see Figure 13 for the prompt) is able to achieve 88.95% accuracy over manually labeled responses, we delineate this limitation for real-world deployment in Section 6.1.1..
The tradeoff between goal reaching and knowledge gaining. Participants diverged in their views on answer-oriented hints helped them answer the question without improving conceptual understanding. Some appreciated strategies that narrowed the response space or partially disclosed the answer (e.g., the hint, ”The name starts with the letter A”, from Figure 5 bottom right interaction), even when these cues offered little additional conceptual knowledge. Others, however, criticized these hints as “misleading” or misaligned with their learning goals. As P-17 reflected, ”[…] A hint which helped me was good but it was not related to the concept itself. I prefer hints which use my external knowledge, but are also somewhat related to the question asked.” This divergent preference reflects a difference in underlying learning objectives (Hoque, 2016; Sönmez, 2017). For participants primarily aiming to maximize quiz performance, hints that directly enabled correct responses were welcomed. For others seeking to deepen their conceptual understanding, context-grounded hints were preferred—even at the cost of leaving the immediate question unanswered.
Unhelpful hints: close similarity to the question or the answer. Participants disliked hints that fell at the opposite ends of the information spectrum, either repeating the questions or revealing the answer directly. Both scenarios limited opportunities for meaningful reasoning. For instance, when asked ”The energy required to remove an electron from a gaseous atom is called?”, one hint rephrased the question (”Think about the process of removing an electron.”), while another gave away the answer (”The term you’re looking for involves ”ionization”.”). Participants described such hints as either redundant or uninformative, reducing both challenge and conceptual engagement. While such hints could support learners focused solely on correctness, as evident by a weak correlation between the self-reported satisfaction ratings and answer leakage ($corr_{pearson}=0.2226$), others found them frustrating, underscoring the importance of hints that guide the learners without shortcutting the cognitive effort.
Hint simplicity and clarity. From a stylistic viewpoint, participants consistently noted that the hints were straightforward and easy to understand. The hints provided clear, focused guidance, allowing learners to quickly grasp the intended strategy or concept.
Figure 5. Interactions between the participants and static (left) and dynamic (right) hint generation strategies. The top interactions depict a successful journey, whereas the bottom interactions illustrate failure to get participants to reach the answer.
5.2. When One Hint Is Not Enough: The Dynamics of Chain-of-hints
While individual hints often provided clear and concise guidance, participants’ responses indicated that a single hint was not always sufficient to successfully answer the question. In many cases, participants required additional hints to fully grasp the concept, resolve ambiguities, or correct misconceptions. Across the 468 questions where hints were requested, only 32.5% were resolved with a single hint, whereas 21.5% required two hints, 13.9% required three, and 32.1% required four. These patterns suggest that learners frequently relied on sequences of hints to fully grasp the concept, resolve ambiguities, or correct misconceptions. Our analysis shows how the chain-of-hints shaped their overall effectiveness, revealing the conditions under which learners benefited from successive guidance as well as the pitfalls that emerged along the way.
Content diversity (or lack thereof). Participants often expected successive hints to provide fresh perspectives rather than reiterations of the same clue. While some participants reported frustration when hints felt repetitive, others valued the consistency, noting that reinforcement sometimes helped clarify the intended direction, as explained by P-26, ”Sometimes, I reached a question where I had no idea what the words even meant - but the hints would help me narrow down my guesses and make much more educated guesses […] I think the hints were good at not providing overlapping information for the most part, and were also quite good and not completely giving away the answer…”. This tension suggests that diversity across a chain-of-hints is not simply about novelty, but about offering complementary information that sustains engagement while gradually deepening understanding.
(a) Distribution of number of attempted answers for static hints.
(b) Distribution of number of attempted answers for dynamic hints.
(c) Average satisfaction ratings relative to the number of attempted answers for static hints.
(d) Average satisfaction ratings relative to the number of attempted answers for dynamic hints.
Figure 6. Distribution of hint chains (top) and corresponding satisfaction ratings for static and dynamic settings across attempted answers.
Coherence and incremental gain. Participants emphasized the value of hint sequences that gradually introduced new information, allowing them to stay challenged while being guided toward the correct answer. Just as important as diversity, many appreciated when hints were coherent (i.e., building logically on one another). This continuity helped them maintain a sense of flow in their learning journey. Such step-by-step scaffolding not only sustained engagement but also fostered a sense of accomplishment as learners progressed. Overall, participants rated hints as informative in roughly 40% of cases at the first position, increasing to about 48% at positions two and three, with a slight decrease to 45% for the fourth position hints, underscoring the role of successive hints in enriching the learning process.
These findings highlight the challenge in designing effective hint chains. In the context of factual question answering, successful hint chains had to avoid being too close to the question to offer little support, or too close to the answer to risk oversimplification, while striking a balance between providing diverse yet relevant information.
5.3. Static vs Dynamic hints
We observed minimal differences in participant performance across the static and dynamic hint conditions, with average success of 7.61/10 and 7.88/10 questions respectively, compared to 6.36/10 questions in the no-hint control setting. Participants requested a total of 639 static hints, and 510 dynamic hints. The perceived difficulty of quiz was also rated similarly across all three sections, suggesting that neither hint strategy provided a decisive advantage at the level of overall outcomes. A Mann-Whitney U test (Mann and Whitney, 1947) revealed no statistically significant difference between the satisfaction ratings ($U=168351,p=0.365$) with negligible effect size (rank-biserial correlation = -0.033). Even after revealing the corresponding hint strategy to participants in the post-quiz survey, opinions remained divided. 20 out of 41 participants found static hints to be more helpful for answering the questions, while 23 out of 41 participants felt dynamic hints better improved their knowledge about the questions. Point-biserial correlation (Kornbrot, 2014) between post-section responses and post-quiz responses revealed that the reported quiz difficulty was not correlated with the post-disclosure preference ($corr=0.067,p=0.675$), indicating that the participants independently rated the quiz difficulty agnostic of hint helpfulness. The post-section hint helpfulness was moderately correlated with the post disclosure informativeness preference obtained from the posttest survey ($corr=0.471,p=0.002$), indicating that participants who found one hint strategy more helpful were likely to maintain their preference even after knowing the hint strategy. However, there is no correlation between the performance difference between static and dynamic, and the post-disclosure preference of answering assistance ($corr=0.063,p=0.694$), indicating that participants’ preferences were not influenced by their actual performance but rather by their subjective experience with the hints. Although neither strategy prevailed overall, participants described clear differences in how they experienced the two hinting approaches.
Static hints–pre-generated in chains of four–were rigid in their progression but often provided diverse facts that enriched the context of the question. For example, ”Reflect on the types of landforms created by magma rising to the surface.”, or ”Research the type of radiation that includes visible light.” (see Figure 5). This pattern was reflected in reported informativeness scores: 39% of static hint chains were labeled informative, compared to only 35% of dynamic chains. However, Chi-squared test (Tallarida and Murray, 1987) revealed no statistically significant difference between the two strategies ($\chi^{2}=0.819,p=0.356$) with no association (Cramer’s V = 0.027). Providing a breadth of information allowed static hints to support sense-making and knowledge-building, but at times the progression felt misaligned with participants’ immediate needs.
Dynamic hints, in contrast, were answer-fixated and more directly oriented toward guiding participants to the correct response. Because these hints were generated adaptively without awareness of the total number of hints in the chain, they lacked the deliberate pacing and planning of the static setting. However, this immediacy was effective when participants lacked sufficient prior knowledge to answer the question. For instance, hints like, ”Think about the rays named after a Greek letter”, or ”Consider a medication that starts with ”a” and is often used for headaches.” (see Figure 5). At the same time, the absence of chain-level planning made it harder to conceal answers: leakage was reported more often in dynamic hints (18%) than in static hints (15%). While Chi-squared test (Tallarida and Murray, 1987) revealed a statistically significance difference between the leakage responses ($\chi^{2}=5.35,p=0.021$), there was no association (Cramer’s V = 0.068). If the participant’s did not engage with the question, (i.e., not attempting any answers) the average satisfaction ratings for dynamic hints was worse than the static hints without engagement. While when they engaged with dynamic hints, their satisfaction ratings were comparable to static hints (see Figure 6 bottom).
Across participants, the overall sentiment remained consistent. Static hints were valued for their contextual richness, while dynamic hints were appreciated for their answer-directed efficiency. Minor exceptions arose due to randomized question distribution, with four participants noting one section to be easier than other, but the broader trend was stable. These findings suggest that static hints may be better suited for fostering understanding and providing transferable knowledge, whereas dynamic hints may be more effective in ensuring task success when learners are struggling.
This contrasting experience highlights a design trade-off between insightfulness and answer-orientation. Further exploration is required to find the right balance, potentially by combining the breadth of static hints with the responsiveness of dynamic hints into a hybrid approach that scaffolds learning while still supporting successful task completion.
5.4. Assessing effectiveness of automatic evaluation metrics
Automatic evaluation metrics are a key step toward scaling hint generation research: they allow us to compare systems without relying exclusively on human annotation. However, their usefulness depends on how well they capture the aspects of hint quality that learners themselves experience. To assess this alignment, we compared a set of automatic evaluation metrics against participant feedback collected in our human study. In total, 41 participants annotated 1,149 hints across 468 questions. Unlike our standardized automatic evaluation setup with a fixed chain size (see Section A), the hint chains here varied in length, providing a more realistic but also more heterogeneous basis for evaluation.
We observed only weak correlations between information gain metrics and participants’ reported informativeness of hints for $InfoGain_{mean}$ ($corr_{pearson}=-0.121$) and $InfoGain_{comb}$ ($corr_{pearson}=-0.142$). These negative but shallow correlations highlight the limitations of relying on LLM-based measures of content quality, which often fail to capture the contextual insightfulness valued by learners. Although our study design did not directly test $Redundancy$ or $Consistency$, we also examined their correlations with user satisfaction. $Redundancy$ showed a weak negative correlation ($corr_{pearson}=-0.245$), while $Consistency$ was essentially uncorrelated ($corr_{pearson}=-0.053$).
For answer leakage, we were able to compare participant reports directly to our automatic metrics. The exact-match metric achieved high precision (0.834) but very low recall (0.075), indicating that it severely underestimated the degree of leakage present. While we implemented $leakage_{EM}$ with a simple string matching approach, that is unable to capture the nuances where hints reveal the answer in plural form (e.g., capillary vs capillaries), or change in part-of-speech (e.g., volcano vs volcanic). We also identified several indirect oversimplifications from hints that remain undetected by a simple syntactic overlap approach. For example, ”The bohr model works only for which atom?”, participants considered the hint ”It is the first element on the periodic table.” to have given away the answer, whereas it was undetected by the $leakage_{EM}$ metric. Conversely, the LLM-based leakage metric offered slightly higher recall (0.183) but suffered from low precision (0.266), indicating that while the learners didn’t consider certain hints to be oversimplifications, the evaluation LLM considered those hint chains to be so, overshooting the answer leakage predictions. For example, for the question ”What is the most abundant metal of the earth’s crust?”, and the provided hints ”Consider which metals are commonly found in rocks and soil.” and ”Think about the primary component of bauxite, a common ore.”, $leakage_{LLM}$ metric considers the hint chain to have leaked the answer (Aluminum), whereas the participant doesn’t indicate that in their responses. These results suggest that neither approach yet provides a reliable estimate of answer leakage from a learner’s perspective.
Taken together, these results indicate that current automatic evaluation metrics only partially reflect learners’ perceptions of hint quality. The weak alignment underscores the need to refine these metrics, either by tuning existing approaches or by developing new measures grounded in interaction data. Our dataset of hint annotations provides a valuable resource for future work toward more learner-centered evaluation.
6. Discussion
Our findings highlight the complexity of automatic hint generation and the limitations of adopting off-the-shelf LLMs as tutoring systems. While participants consistently valued certain hint characteristics such as simplicity, novel knowledge perspectives, content diversity, and incremental scaffolding, their preferences diverged on various other aspects. For instance, some learners favored answer-focused hints that guided them toward the answer, while others found such feedback unhelpful in improving their learning. Beyond feedback on the hints generated by our system, participants also emphasized some unmet needs, with the hints unable to help bridge the linguistic barriers or connect well to their prior knowledge. Below, we discuss design considerations for advancing hint generation beyond factual question answering (Section 6.1), and we position this research problem as a potential avenue for addressing the growing challenge of cognitive offloading among learners (Section 6.2).
6.1. Design Implications Beyond Factual Hinting
The revised Bloom’s taxonomy of educational goals (Bloom et al., 1956; Anderson and Krathwohl, 2001) organizes cognitive processes from remember and understand through apply, analyze, evaluate, and create, while distinguishing knowledge goals between factual, conceptual, procedural, and metacognitive dimensions. Our study deliberately focused on factual knowledge at the lower levels of this hierarchy, where hints primarily supported remembering and understanding. This scope allowed us to systematically examine the role of hinting, but also revealed several limitations in current systems, for instance, challenges in robust answer assessment and misalignment between automatic evaluation metrics and learner perceptions. As one moves upward in Bloom’s hierarchy, such limitations become even more consequential: supporting conceptual reasoning, procedural mastery, or metacognitive reflection requires richer forms of adaptation, sensitivity to learner goals, and integration of affective and self-regulatory scaffolding. Next, we outline design implications that extend beyond factual hinting, drawing on our findings to point toward future directions for building tutoring systems capable of supporting higher-order learning.
6.1.1. Building Trustworthy Foundations
Robust Answer Assessment. To evaluate the reliability of our answer assessment module for the user study, we manually annotated 697 attempted answers from a pilot study. The in-context learning system achieved an accuracy of 88.95% in correctly labeling learner responses. While answer assessment was not the central focus of this work, we considered this level of performance sufficient for conducting a controlled user study. Nonetheless, even in the relatively constrained setting of short-form factual questions, our system encountered challenges in reliably assessing some learner responses. For example, while it correctly accepted “CO2” as an equivalent to “carbon dioxide,” it failed on other equivalent forms such as “OH-” for hydroxide ions, and in one case incorrectly marked “ions” as fully correct. Such inconsistencies highlight the fragility of off-the-shelf LLM-based assessment pipelines and underscore how errors, if unaddressed, can affect learner trust and the validity of downstream evaluation.
The difficulty compounds when moving beyond factual recall, as assessment approaches vary across domains. In subjects such as mathematics, robust evaluation requires formal verification of solutions and scrutiny of intermediate steps to detect logical errors (Sangwin and Köcher, 2016; Lan et al., 2015; Beevers and Paterson, 2003). In contrast, domains like history or biology demand capabilities such as argumentation mining and semantic overlap analysis (Lippi and Torroni, 2016; Lawrence and Reed, 2020; Li et al., 2025a). For instance, in answering the open-ended question “Why did the Western Roman Empire fall?,” there could be multiple possible responses, but an effective assessment would align the learner’s reasoning with key factors outlined in reference solutions. Building reliable assessment modules that are rigorously tested and aligned with the cognitive goals of instruction is therefore both a significant design challenge and a necessary foundation for developing effective, equitable feedback systems that support higher-order learning objectives.
Automatic Evaluation of Hint Quality. Our study also highlights the limitations of existing automatic evaluation metrics for hint generation. While such metrics are essential for scaling research without relying exclusively on human annotation, our analysis revealed that they are, at best, weakly aligned with learners’ perceptions of hint quality. For example, information gain metrics showed shallow negative correlations with participant ratings of informativeness, suggesting that LLM–based measures fail to capture the contextual insightfulness learners value. Similarly, our automatic leakage metrics exhibited trade-off. Exact-match rules achieved high precision but severely underestimated leakage, while LLM-based approaches captured more cases but at the cost of low precision. Redundancy and consistency measures also showed little to no correlation with learner satisfaction. These findings suggest that current metrics do not yet provide a reliable proxy for learner experience. Progress in this space will require either refining existing metrics or developing new ones grounded in interaction data and learner-centered perspectives. Our annotated dataset of 1100+ hints across 450+ questions could support development of more trustworthy evaluation frameworks.
LLM-as-a-Judge. Using LLMs as evaluators of natural language output-commonly referred to as LLM-as-a-judge-has gained traction in NLP research (Gu et al., 2024). Compared to resource-intensive human studies, it offers easy setup, low expense, and circumvents challenges such as IRB approvals, funding constraints, and participant recruitment. Yet this approach comes with significant limitations. Prior work shows that LLM-based evaluators are prone to systematic biases (Chen et al., 2024a; Ye et al., 2024; Shi et al., 2024; Wang et al., 2023a) and are vulnerable to adversarial manipulation (Tong et al., 2025). Self-assessment studies, such as Krumdick et al. (2025), demonstrate that while LLMs may show high overall agreement with human annotators, they often fail on subsets of tasks they cannot themselves solve. Similarly, Szymanski et al. (2025) found that LLM-based evaluation performs poorly in expert domains (e.g., dietetics, mental health), with substantially lower agreement than human judges. Beyond performance metrics, qualitative investigations raise broader concerns: Kapania et al. (2025) showed that when LLMs are used to simulate participants, their outputs often lack contextual depth, obscure marginalized perspectives, and amplify researcher positionality. There has been recent interest in using LLM-as-a-judge for evaluating education technologies or simulating student behavior (Dinucu-Jianu et al., 2025; Li et al., 2025b; Benedetto et al., 2024; Lu and Wang, 2024; Mannekote et al., 2025; Yuan et al., 2025; Zhang et al., 2024). However, educational evaluation is uniquely sensitive, as it involves assessing nuanced learner behaviors, metacognitive processes, and affective states—dimensions that the use of LLMs might oversimplify or misrepresent. Without stronger validation and safeguards, we advocate caution against relying on LLMs as judges to drive the design of tutoring systems, as premature adoption risks misaligning feedback with learners’ needs, and undermining trust in educational technologies.
6.1.2. Designing for Learner Diversity and Goals
Designing for diverse learners. Our study highlights the limitations of relying solely on off-the-shelf general purpose models in educational context. Recent years have seen a shift from specialized models towards large, general purpose foundational models (Bommasani, 2021; Nori et al., 2023; Li et al., 2024). While training on massive datasets has produced highly capable systems with emergent abilities (Wei et al., 2022; Berti et al., 2025), this trend risks discarding lessons from decades of research in the ITS community, which emphasized specialized, learner-centric frameworks (VanLehn, 2011; Zhang and Jia, 2017; Conati et al., 2021). Our findings also show that learners have different language proficiencies, educational backgrounds, and goals, which shape how they interpret and act on hints. As one participant (P-22) explained, ”…I know how to answer them in my native language, but in English I get limited by my vocabulary. (I used to be able to spell and handwrite these words, but it’s already been five years since I graduated from high school! I’ve forgotten how to write them, though I can still recognize the words when I see them.)” P-22’s experience exemplifies the importance of designing systems that accommodate linguistic diversity and recognize that learners’ needs vary widely. We argue for a balanced approach that combines the reasoning capabilities of LLMs in conjunction with insights from specialized models to better cater to individual learner needs, while adopting non-invasive methods for modeling preferences. One promising direction is to infer learner goals and prior knowledge from interaction data. For example, Liu et al. (2022), utilizes long short-term memory cells (Hochreiter and Schmidhuber, 1997) for knowledge estimation (Corbett and Anderson, 1994) of student’s current understanding based on their past responses for open-ended program synthesis. We can adopt user behavior modeling techniques from the vast literature of recommendation systems (He et al., 2023) to measure progress and adapt the hints to the learner’s preferences. Beyond personalization, future work should also rigorously evaluate tutoring approaches across more diverse populations to ensure that adaptive systems provide equitable and inclusive support.
Aligning with learner goals. Every learner brings different objectives depending on the task and context (Hoque, 2016; Sönmez, 2017). Our findings reveal a similar learners’ preference variability, where some participants valued question-agnostic hints that guided them toward answers, while others who sought deeper understanding found these unhelpful. While providing context agnostic feedback to facilitate successful responses from learners improves short-term satisfaction, it hinders in development of critical thinking. While it is important to identify individual learning goals to provide the most suitable feedback, further caution is required to prevent over-reliance and over simplification. In Section 6.2, we discuss this tension of satisfaction with cognitive offloading in depth. Besides that, even identifying these learning goals isn’t a straightforward task, as prior work has shown, learners often lack the metacognitive awareness to articulate their own goals (Karpicke et al., 2009; Ku and Ho, 2010), and their learning preferences evolve over time (Gurpinar et al., 2011; Andrews and Tynan, 2011). Thus, tutoring systems must balance explicit input from learners with adaptive mechanisms that infer goals through interaction patterns. Metrics such as satisfaction ratings or longitudinal engagement may provide richer signals for tailoring support, pointing to the need for nuanced feedback frameworks that remain grounded in pedagogical values.
6.1.3. Beyond Cognition: Toward Holistic Learning Support
Need for mixed-initiative support. Prior work in intelligent tutoring shows that learners often delay or avoid seeking hints (Aleven and Koedinger, 2000), whereas human teachers frequently provide just-in-time guidance to prevent disengagement (Novak et al., 1999). Moreover, research has shown that low-performing students may lack the metacognitive skills necessary to make effective pedagogical decisions, such as when to request support (Aleven and Koedinger, 2000; Roll et al., 2014). Addressing this challenge, Ju et al. (2022) introduced a mixed-initiative tutoring framework using reinforcement learning, which successfully boosted the performance of low-achieving learners in a pretest–posttest study. While these results are promising, designing AI systems that can emulate the natural balance between learner autonomy and timely support remains a critical open challenge for future research.
Toward scaffolding self-regulated learning. While mixed-initiative frameworks can compensate for learners who struggle to recognize when support is needed, cultivating metacognitive skills themselves is essential for fostering critical thinking (Ku and Ho, 2010). Tutoring systems should therefore not only deliver assistance but also nurture learners’ capacity for self-regulation. Two pedagogical paradigms with strong empirical support offer promising directions: self-regulated learning (SRL) (Zimmerman, 1990, 2002; Prasad and Sane, 2024) and learning-by-teaching (LBT) (Frager and Stern, 1970). Both approaches have long-standing evidence of effectiveness in enhancing metacognition and deep learning (Allen and Feldman, 1973; Bargh and Schul, 1980; Chi et al., 2001; Cohen et al., 1982; Cheng, 2011; Bouchet et al., 2012), and recent advances in AI have opened new possibilities for embedding them in digital environments. For instance, Ge et al. (2025) demonstrate that LLM-assisted SRL can significantly improve regulation skills when scaffolded within a gamified environment. Likewise, Rogers et al. (2025) show how positioning LLMs as “tutees” enables learners to practice learning by teaching, thereby strengthening conceptual understanding as well as metacognitive reflection. Building on these insights, we envision future hint generation systems that extend beyond factual recall: systems that scaffold reflection, gradually fade support, and empower learners to assume increasing responsibility for their own learning. Such designs not only reduce risks of over-reliance but also promote the development of independent, critical thinkers.
6.2. Over-reliance and Cognitive Offloading
Recent years have seen rapid adoption of AI tools across everyday life including education, where students increasingly turn to AI assistance for help with assignments, writing, and problem-solving (Pitts et al., 2025a, b; Kanont et al., 2024; Lin, 2024). Prior work has documented this growing dependence, noting that learners often use these systems not only as aids but also as substitutes for their own cognitive effort (Adiyono et al., 2025). While such tools can provide immediate support, they also raise pressing concerns about their long-term effects on learning. Our findings echo this trend, as many participants preferring low-effort, context-agnostic hints that directly facilitated answering questions. This underscores a critical tension between short-term satisfaction and the risk of accumulating long-term cognitive debt. We contend that carefully designed hint generation systems could support learning without fostering over-reliance.
Several studies offer evidence that over-reliance on AI assistance may negatively affect cognitive engagement. Kosmyna et al. (2025) used electroencephalography (EEG) to measure neural activity during essay writing, comparing students who wrote independently against those who used LLM-based support. They found that AI-assisted writers exhibited markedly weaker neural connectivity patterns, suggesting reduced cognitive coupling during the task. Similarly, Gerlich (2025), through a mixed-method approach, found that younger learners reported higher dependence on AI tools and demonstrated lower critical thinking scores compared to older participants. Importantly, their findings also show that educational attainment correlates with stronger critical thinking skills regardless of AI use, pointing to both risks and potential mitigators of over-reliance. While these technologies can enhance writing proficiency and boost confidence, they also introduce risks of diminished creativity, over-reliance, and ethical concerns like plagiarism and systemic biases (Zhai et al., 2024).
Together, these findings highlight the urgency of designing educational technologies that support learners without undermining cognitive engagement. Rather than fully outsourcing intellectual effort, tutoring systems should aim to preserve opportunities for reflection, reasoning, and persistence. Recent research has taken steps toward this goal—for example, tools that emphasize explainability (Hoq et al., 2025), promote metacognitive reflection (Kumar et al., 2024; Yuan and Hu, 2024), or deliberately restrict direct answer-giving (Gupta et al., 2025). These approaches seek to balance the benefits of AI assistance with the preservation of deeper thinking.
Hint generation represents one such framework. By design, hints avoid revealing the complete answer, instead scaffolding learners’ problem-solving processes while encouraging persistence and self-efficacy. Learners often seek AI assistance when they are stuck, and unlike readily available AI systems that may provide full solutions, hint generation frameworks emphasize gradual support grounded in pedagogy principles, enabling learners to build knowledge rather than bypass it. Prior work in intelligent tutoring consistently shows that effective hints can increase motivation, improve self-efficacy, and foster a sense of accomplishment (Rogers et al., 2025; Mohammadi et al., 2025; VanLehn, 2011; Ma et al., 2014). Our findings reinforce this perspective: participants generally rated hints as helpful for answering questions, with several explicitly noting their value in promoting deeper thinking. For instance, P-22 described hints as ”hints encouraged to make associations and engage in reflection from a scientific perspective,” P-37 noted they ”encouraged active thinking,” and P-40 highlighted they ”encouraged me to look for other ways to say things.” Taken together with prior research, these results suggest that automatically generated hints can serve as a middle ground: providing support without fostering harmful dependence, and helping learners sustain their cognitive engagement in the presence of AI.
7. Limitations
Our study has several limitations that should be acknowledged when interpreting the findings. First, there are recruitment-related constraints. While our target demographic was high school students, ethical considerations around exposing minors to early-stage AI technology led us to recruit recent high school graduates instead. As a result, our participant pool consisted primarily of university students.
Second, our study design reflects several methodological choices that constrain the scope of our findings. By focusing on factual questions, we primarily probed for remembering and understanding, rather than deeper levels of cognition such as application or analysis. This limits the generalizability of our insights to higher-order learning tasks. Moreover, our learner-driven setup assumes that participants possess sufficient metacognitive awareness to judge when to request hints. Prior research (Karpicke et al., 2009; Ku and Ho, 2010) suggests that this assumption does not always hold true, as many learners lack these skills, which in turn restricts the broader applicability of our proposed hint generation system. Additionally, our study also doesn’t probe the participants’ views of AI technologies, a factor that can affect the applicability of the results to broader learner population.
Finally, there are limitations tied to the implementation of our hinting strategies. The dynamic hinting condition was unaware of the hint chain length, making it more difficult to plan hint sequencing and leading to a somewhat uneven comparison with the static condition. Furthermore, our system was agnostic to learners’ backgrounds and prior knowledge: while hints adapted superficially to student responses, they did not provide deeper personalization that could account for individual learning histories or subject expertise. Besides prior knowledge, our design also didn’t take into consideration affective dimensions like frustration, disengagement, or overconfidence, even though they are critical for learning (Mao and Li, 2010; Cunha-Pérez et al., 2018; Yadegaridehkordi et al., 2019; Hasan et al., 2020).
Taken together, these limitations highlight important opportunities for future work. Recruiting participants closer to the intended learner demographic, extending evaluation to higher-order learning tasks, and developing adaptive systems that better account for learner diversity are all critical next steps. Addressing these gaps will help refine automatic hint generation into a more robust and equitable tool for supporting diverse educational contexts.
8. Conclusion
This paper examined the capabilities of large language models to generate and evaluate hints for science education. We benchmarked 18 open-source models using five automatic evaluation metrics grounded in learning sciences, each designed to capture a distinct attribute of effective feedback. We investigated two complementary hinting strategies: static hints, pre-generated for each problem, and dynamic hints, adapted to learners’ progress. Our quantitative study with 41 participants underscores both the promise and complexity of providing automated hints. Learners valued certain attributes that made hints effective, while also identifying aspects that hindered their learning experience. Building on these insights, we propose design principles for future systems that align with Bloom’s taxonomy of educational goals and emphasize learner-centered design. More broadly, we position automatic hint generation as a key research direction to address the growing challenge of cognitive debt among next-generation learners.
Appendix A Automatic Evaluation of Scientific Chain-of-Hints
We describe and motivate each automatic evaluation metric in Section A.1, and discuss how capable various open-source LLMs are in generating scientific chain-of-hints in Section A.2.
A.1. Automatic Evaluation Metrics
Figure 7. Information gain of ROUGE-L on validation set for randomly selected 30 hints using Gemma-2, Llama-3, Mistral, Qwen2.5 and Qwen3 models.
Information Gain. To measure the extent to which a hint can add information to the knowledge base of a user to help answer a question is one of the most important attributes of a hint. To interpret whether a user will be able to gain the necessary information from a hint, we use a large language model to measure the information gain from a hint. More specifically, we measure the difference in performance of the evaluation model with and without the hints. Due to the tendency of LLMs to generate verbose responses, we use ROUGE-L recall (Lin, 2004) to match the generated response to the correct answer as a scalable approximation for assessment. We propose two different variants for this metric: i) average information gain across all hints ($InfoGain_{mean}$), and ii) combined information gain from all hints ($InfoGain_{comb}$), defined as follows -
(1) $\begin{split}InfoGain_{mean}(\mathbf{h})&=\frac{1}{k}\sum_{i=1}^{k}ROUGE_{recall}^{L}\bigl(LLM(q,h_{i}),a\bigr)\\ &\quad-ROUGE_{recall}^{L}\bigl(LLM_{QA}(q),a\bigr)\end{split}$
(2) $\begin{split}InfoGain_{comb}(\mathbf{h})&=ROUGE_{recall}^{L}\bigl(LLM(q,\mathbf{h}),a\bigr)\\ &\quad-ROUGE_{recall}^{L}\bigl(LLM_{QA}(q),a\bigr)\end{split}$
where h is a chain-of-hints corresponding to question $q$ and answer $a$, and $LLM_{QA}(.)$ is the question answering LLM.
[…]
Redundancy. Variation theory suggests that contrasting examples and perspectives help learners discern key features of a concept (Ling Lo, 2012). Lomibao and Ombay (2017)’s findings aligned with the theoretical argument of variation theory, the students exposed to repetition with variation to achieved significantly higher performance and improved knowledge retention in mathematical problem solving. To evaluate the diversity of perspectives in a chain-of-hints, we measure its self-referenced redundancy using the Sentence-BERT embeddings (Reimers and Gurevych, 2019) as proposed by Chen et al. (2021a) -
(3) $Redundancy(\textbf{h})=\sum_{i=1}^{k}max_{j:i\neq j}Cos_{sim}(h_{i},h_{j})$
where $j:i\neq j$ denotes that we don’t consider similarity of hint j ($\tilde{h_{j}}$) with itself.
Consistency. To ensure the high quality of information within the hints, we measure the consistency of the hints using AlignScore (Zha et al., 2023), a factual consistency evaluation metric that uses information alignment between context and claims to measure the factual consistency.
(4) $Consistency(\textbf{h})=AlignScore(context,context)$
Leakage. Learners construct cognitive representations when they are guided to derive the solution themselves, molding them into deep conceptual thinkers with the ability of question more, seeking to understand rather than to only memorize (Rillero, 2016). However, revealing the correct answer prematurely disrupts this cognitive engagement, hindering their learning journey by oversimplifying the problem (Brown et al., 2014). To capture this, we adopt two metrics to evaluate answer leakage in generated hints - 1) Exact Match (EM), where answer is leaked if the exact answer string is present in the hint, and 2) LLM prompt (LLM), where we use the large language model’s in-context learning capabilities to identify presence of answers in the hints, capturing more diverse types of leakages like paraphrased answers (e.g., ”volcanic activity” instead of ”volcanoes”) or abbreviations (e.g., ”CO2” instead of ”carbon dioxide”). We use Gemma3-27B model for the Leakage (LLM) evaluation metric (see Figure 12 for the exact prompt used by this metric).
Readability. Hints should be expressed in a clear and simple language to ensure that the learners can easily grasp the underlying concept. Avoiding unnecessary complexity in conveying the critical problem solving strategy enhances the usefulness of hints, reinforcing the well-established principle of direct instructions (Kozloff et al., 1999; Kim and Axelrod, 2005; Rosenshine, 2008). To grasp this linguistic clarity in the generated hints, we use three different readability measures - i) Flesch-Kincaid (FK) grade level readability score (Kincaid et al., 1975), ii) Flesch reading ease (FRE) (Flesch, 1979), and iii) Dale Chall (DC) readability score (Dale and Chall, 1948). The Flesch-Kincaid grade level score and Flesch reading ease are scored using a linear combination of number of words per sentence and number of syllables per word, whereas the Dale Chall readability score utilizes a list of 769 words that 80% of fourth-grade students are familiar with to provide a readability score. We implement these readability scores using py-readability-metrics [^5] https://github.com/cdimascio/py-readability-metrics.
Aggregate Score. In order to rank the baselines and finalize the best model for our human evaluation study, we formulate a simple aggregate score over the automatic evaluation metrics [^6] We don’t use readability for the aggregate function as all the hints generated by the baselines have similar readability scores, and generate well-formulated sentences. as follows -
$Aggregate(\mathbf{h})=\tfrac{1}{4}\left\{\begin{split}&InfoGain_{comb}(\mathbf{h})+Consistency(\mathbf{h})\\ &+\bigl(1-Redundancy(\mathbf{h})\bigr)+\bigl(1-leakage_{EM}(\mathbf{h})\bigr)\end{split}\right\}$
Table 3. Automatic evaluation metric results for static hint generation setting.
Baseline $InfoGain_{mean}\uparrow$ $InfoGain_{comb}\uparrow$ $Redundancy\downarrow$ $Consistency\uparrow$ $Readability_{DC}\downarrow$ $Leakage_{EM}\downarrow$ $Leakage_{LLM}\downarrow$ $Aggregate\uparrow$
DeepSeek-R1-1.5b 0.214 0.421 0.625 0.199 10.810 0.369 0.859 0.406
DeepSeek-R1-7b 0.206 0.434 0.641 0.334 10.409 0.220 0.543 0.476
DeepSeek-R1-8b 0.161 0.418 0.572 0.302 10.152 0.070 0.557 0.520
DeepSeek-R1-14b 0.121 0.423 0.563 0.396 9.514 0.045 0.449 0.553
DeepSeek-R1-32b 0.148 0.450 0.581 0.420 10.183 0.087 0.472 0.550
Gemma3-1b 0.119 0.369 0.529 0.298 10.929 0.053 0.170 0.521
Gemma3-4b 0.122 0.419 0.572 0.395 10.139 0.054 0.417 0.547
Gemma3-12b 0.047 0.393 0.524 0.382 9.959 0.019 0.337 0.558
Gemma3-27b 0.028 0.354 0.539 0.332 10.070 0.011 0.479 0.534
Qwen3-0.6b 0.234 0.513 0.741 0.450 10.173 0.573 0.972 0.412
Qwen3-1.7b 0.198 0.484 0.681 0.356 10.163 0.329 0.792 0.458
Qwen3-4b 0.190 0.470 0.634 0.383 9.438 0.232 0.729 0.497
Qwen3-8b 0.168 0.473 0.613 0.393 9.508 0.178 0.628 0.519
Qwen3-14b 0.126 0.416 0.552 0.307 9.455 0.043 0.491 0.532
Qwen3-30b 0.131 0.414 0.559 0.341 9.905 0.024 0.463 0.543
Qwen3-32b 0.138 0.436 0.577 0.354 9.454 0.034 0.664 0.544
Mistral-24b 0.103 0.429 0.544 0.408 9.558 0.022 0.382 0.568
Phi4-14b 0.191 0.441 0.592 0.337 10.555 0.046 0.530 0.535
Table 4. Automatic evaluation metric results for dynamic hint generation setting.
Baseline $InfoGain_{mean}\uparrow$ $InfoGain_{comb}\uparrow$ $Redundancy\downarrow$ $Consistency\uparrow$ $Readability_{DC}\downarrow$ $Leakage_{EM}\downarrow$ $Leakage_{LLM}\downarrow$ $Aggregate\uparrow$
DeepSeek-R1-1.5b 0.358 0.504 0.833 0.388 10.439 0.723 0.998 0.334
DeepSeek-R1-7b 0.371 0.508 0.891 0.463 10.555 0.707 0.998 0.343
DeepSeek-R1-8b 0.172 0.425 0.713 0.256 10.480 0.121 0.847 0.462
DeepSeek-R1-14b 0.201 0.447 0.777 0.375 10.067 0.172 0.876 0.468
DeepSeek-R1-32b 0.251 0.467 0.769 0.364 10.133 0.241 0.888 0.455
Gemma3-1b 0.205 0.395 0.744 0.367 10.995 0.165 0.517 0.464
Gemma3-4b 0.134 0.392 0.716 0.392 10.471 0.064 0.303 0.501
Gemma3-12b 0.093 0.413 0.648 0.336 10.097 0.063 0.245 0.509
Gemma3-27b 0.109 0.402 0.670 0.271 10.159 0.041 0.316 0.491
Qwen3-0.6b 0.270 0.504 0.975 0.491 10.881 0.872 0.987 0.287
Qwen3-1.7b 0.372 0.507 0.925 0.513 10.330 0.743 0.987 0.338
Qwen3-4b 0.340 0.505 0.890 0.426 10.230 0.493 0.979 0.387
Qwen3-8b 0.223 0.487 0.828 0.439 10.507 0.272 0.861 0.457
Qwen3-14b 0.268 0.469 0.774 0.366 10.557 0.087 0.892 0.493
Qwen3-30b 0.200 0.453 0.721 0.365 10.077 0.051 0.641 0.512
Qwen3-32b 0.263 0.472 0.791 0.362 10.614 0.095 0.955 0.487
Mistral-24b 0.158 0.437 0.665 0.384 9.581 0.036 0.357 0.530
Phi4-14b 0.426 0.501 0.811 0.186 10.861 0.523 0.898 0.338
A.2. Automatic Evaluation Results
We present the performance of all 18 baseline models for static and dynamic hint generation at Table 4 and 4 respectively. In this section, we discuss our observations from this automatic evaluation process, and finalize a baseline for the subsequent human evaluation user study.
Figure 8. Pearson’s correlation matrix of automatic evaluation metrics for static (left) and dynamic (right) hint generation.
Trends across evaluation metrics. We compute the Pearson’s correlation across the automatic evaluation metrics to identify several trends over the baselines (refer to Figure 8 for the correlation matrices). We observe that the variants of evaluation metrics are highly correlated with each other. E.g., for $InfoGain_{mean}$ and $InfoGain_{comb}$ is 0.77, $Readability_{FRE}$ and $Readability_{FK}$ is -0.94 for static hint generation. $InfoGain$ and $Leakage$ are also heavily correlated ($corr>0.7$), understandably so, since the baselines with tendency to leak the answers oversimplifies the problems for the $InfoGain$ evaluator LM. This is further evident from the pareto-optimal curves between the two metrics (refer to Figure 9 top-left and top-right), with a diverse range of models at the pareto-front, with the smaller models with high leakage or low information gain at the extreme ends, and optimal baselines at the elbow of the curve. We also observe redundancy to be highly correlated with both information gain and leakage ($corr>0.8$), likely because the baselines that tend to leak answer lack semantic diversity in their hints, leading to high redundancy along with higher information gain.
Figure 9. Pareto-optimal fronts for static hint generation.
Although $Consistency$ is also a semantic evaluation metric, it doesn’t correlate well with other semantic metrics, except $InfoGain_{comb}$ ($corr=0.48$) in static setting. We posit this to be likely because $Consistency$ is the only reference-based evaluation metric, aiming to evaluate the factual correctness of the generated hints compared to the context. As $InfoGain_{comb}$ indirectly measures the semantic information captured by a chain-of-hints towards helping answer the question, the moderate correlation between the two metrics is justified. For dynamic setting on the other hand, $Consistency$ is correlated with RedundancyRedundancy ($corr=0.63$) and $Leakage_{EM}$ ($corr=0.49$) as well, most likely because of significantly higher leakage rates and redundant hints compared to the static hint generation setting.
We observe limited variability across all three variants of the readability scores, with the results roughly interpretable to be between high school (8.0-8.9) to college (9.0+) level according to the Dale-Chall (high school: 8.0-8.9, college: 9.0-9.9), Flesch reading ease (high-school: 60-50, college: 50-30), and Flesch-Kincaid (high school: 9-12, college: 12-15). Although all readability results follow similar interpretation, we observe that $Readbility_{DC}$ is moderately correlated to the other two metric, likely due to its contrasting strategy of using the ratio of difficult words instead. Being a stylistic measure, readability scores don’t correlate well with the other semantic metrics.
Trends across static and dynamic hints We observe some consistent differences in scores between the static and dynamic hints across all baselines. Static hints consistently outperformed dynamic hints $($$\Delta Aggregate=-0.078$)). This was because dynamic hints performed worse than static hints in all evaluation metrics except readability. Compared to static hint generation, the dynamic hints had increased redundancy ($\Delta Redundancy=0.195$), increased answer leakage ($\Delta Leakage_{EM}=0.169,\Delta Leakage_{LLM}=0.201$), leading to slightly increased information gain ($\Delta InfoGain_{ind}=0.098$, $\Delta InfoGain_{comb}=0.029$), and marginal change in consistency ($\Delta Consistency=0.019$). We posit the poor quality of distractors used to steer the generation of dynamic hints to be the leading cause of this worsened performance. For example, for the question ”What kind of waves are sound waves?”, the dataset provides the distractors ”spinning”, ”external”, and ”internal” against the correct answer ”mechanical”. Providing pseudo relevant attempted answers forced the baselines to either repeat their point across multiple hints, or reveal the answers directly or indirectly. While using these distractors as attempted might not perfectly represent a learner’s attempts, they still act as a useful tool to identify the robustness of model response across different settings, helping us identify models capable of providing useful feedback for real-world adoptability.
Trends across model size. To formulate guidelines on what models to use for future research, we explore trends over three model families with multiple variants to observe how their hint generation capabilities varies across scale. We present the results for the largest family of models Qwen3 (Yang et al., 2025) in Figure LABEL:fig:scale_autoeval_main, and results for other two model families DeepSeek-R1 (Guo et al., 2025) and Gemma3 (Team et al., 2025) in Figure 14. We observe that the aggregate score increases with the model size, with static hints outperforming dynamic hints. This gap in performance decreases with the increase in number of parameters, indicating that the larger models are more capable at generating adaptive hints. The biggest contributor to this decrease in performance can be attributed to the answer leakage, where smaller models (<<4b parameters) have the exact answer string for over 20% of instances for Qwen3 and DeepSeek-R1 models. This leakage oversimplifies the problem, leading to an observable difference in the $InfoGain_{mean}$ metric, which measures the average information gain across a chain-of-hints. This effect is not as evident in the $InfoGain_{comb}$ metric, likely because it more holistically gauges the information gain of the entire chain, which even for larger models with less leakage, provide ample information to help the information gain evaluator LM. Smaller models are also not very expressive, with more redundant hint sequences compared to larger models, that provide more diverse perspectives due to their better memorization and reasoning capabilities. While there’s a definitive trend of improved hint quality with an increase in size, for a simpler task of factual scientific question answering, the performance plateaus after a certain size, with models between 12-14b parameters performing sufficiently well off-the-shelf. For a resource constrained domain of education, we believe further finetuning these medium size models to be the most pragmatic next direction.
Appendix B Prompts
We provide the prompts we used throughout our study in this section. Figures 10 and 11 describe the static and dynamic hint generation prompts we used for both automatic evaluation benchmarking and the human evaluation study. Figure 12 contains the prompt we used to measure the LLM-based answer leakage in $leakage_{LLM}$ evaluation metric (using Gemma3-27B model (Team et al., 2025)). Figure 13 comprises of the answer assessment prompt we used to resolve fuzzy matching of participants’ attempted answer with the ground truth answer (using OpenAI’s o4-mini model). The answer assessment was preceded by an exact string match to increase the assessment performance due to the non-deterministic nature of LLM responses. While the figures present the prompt in two parts, a system prompt and the user prompt, in practice, we treated them as a single string to ensure a fair comparison across all models. We leave the finetuning of these prompts to future works.
Figure 10. Prompt used to generate static hints.
Figure 11. Prompt used to generate dynamic hints.
Figure 12. Prompt used for $Leakage_{LLM}$ evaluation metric.
Figure 13. Prompt used for answer assessment in our user interface backend.
Figure 14. Performance of Qwen3 (top), DeepSeek-R1 (middle), and Gemma3 (bottom) model families using automatic evaluation metrics. Orange (solid) bars denote the static hint generation results and blue (striped) bars denote the dynamic hint generation results.